OpenStudy (anonymous):

Need answers to this question on Harmonic oscillator, so that i can compare your solution to mine. A homogeneous stick of length l = 1.0 m and mass m = 0.30 kg is hinged at one end and is connected to a spring with k = 40 N/m at its other end. The stick performs a harmonic oscillation with small deflections. Friction can be neglected. The moment of inertia of a stick around one end is J= (M.I^2)/3 a) Determine the equation of motion for the angular oscillation of the stick. b) Calculate its oscillation period. @OCW Scholar -

OpenStudy (anonymous):

What is the orientation of the setup? How should we account for gravity?

OpenStudy (anonymous):

I am not fully sure about this .. I confirm this in a few days 1/2(Iw^2)=(kx^2)/2 replacing I i got w= 4.5 rad/s